# How to Complete Domain Tuning while Keeping General Ability in LLM: Adaptive Layer-wise and Element-wise Regularization

### v1

@@ 5 Experiments
5.1 Backbone LLM
We evaluate the proposed method using two mainstream LLM: (1) GPT-J (Wang & Komatsuzaki, 2021) is a 6-billion-parameter transformer-based model developed by EleutherAI. It is pretrained on a diverse corpus of texts, making it suitable for a wide range of natural language understanding and generation tasks. (2) LLaMA-3 (Dubey et al., 2024) is the second-generation open-source language model developed by Meta AI. It is designed with enhanced efficiency and scalability, offering state-of-the-art performance across various benchmarks. These models vary in architecture and parameter count, enabling a robust evaluation of our method.
5.2 Tasks, Metrics and Hyperparameters
$\mu$ Task (Domain Ability): The $\mu$ task evaluates the ability to adapt to specific tasks while mitigating catastrophic forgetting of general knowledge. We select three representative tasks: (1) Medical task: Using the MedMCQA dataset (Pal et al., 2022) related to medical information extraction and understanding. (2) Scientific task: Using the SciQ dataset (Welbl et al., 2017) related to scientific literature and reasoning. (3) Physics task: Using the PiQA dataset (Bisk et al., 2020) related to physical reasoning and problem-solving.
@@ 5.2 Tasks, Metrics and Hyperparameters
5.3 Baseline Methods
@@ 6 Results and Analysis
6.1 Accuracy Comparision on General and Domain
As shown in Table 1, our method achieves better preservation of general ability (as reflected by the lowest PPL) while maintaining domain-specific accuracy comparable to, or even better than, previous methods. This demonstrates that our approach effectively balances maintaining high domain accuracy with minimizing the loss in general ability, setting a new standard in preserving both aspects simultaneously.
@@ 6.2 Complexity Comparision
We compare our ALoRA method with the previous SOTA method, EWCLoRA, from two aspects: the time required for importance calculation and the storage memory needed. The experimental results, as shown in the figure, show that our method is nearly 20 times faster and requires only 10$\sim$15of the storage memory compared to the previous method, demonstrating the practicality of our approach.
Time Complexity: The experiments were conducted on an A800 GPU to evaluate the time complexity of our method in comparison with EWCLoRA. For EWCLoRA, the Fisher matrix computation followed the approach described in the original paper, using 20,000 randomly sampled data points from the Pile dataset with a maximum batch size of 8. In contrast, for our method, the time measurement was based on 5 training epochs, a setting determined through empirical evaluation to achieve optimal performance. The experimental results show that for GPT-J-6B and LLaMA-3-3B, EWCLoRA requires 27.17 and 25.97 hours, respectively, to compute the importance matrix, while our ALoRA method only takes 1.15 and 1.19 hours.
@@ 6.3 Regularization Coefficient Analysis
Figure 6 demonstrate the effect of the regularization coefficient $\varphi$ in Equation 18 on PPL and accuracy across three tasks. As $\varphi$ increases, PPL gradually decreases, indicating a stronger emphasis on preserving general ability. Higher values of $\varphi$ correspond to better general ability retention. However, as shown in Figure 6b, increasing $\varphi$ negatively impacts the average accuracy on PiQA. Based on these observations, we select $e^{-3}$ as the optimal value for the regularization coefficient, as it strikes a balance between task performance and general ability (lower PPL).
@@ 6.4 Parameters Importance Visualization
Figure 7 highlights the importance in Equation 15 of q_proj and v_proj layers for general-purpose capabilities during the LoRA fine-tuning process. The heatmap illustrates that the v_proj layers, particularly in the first four and the last layer, are crucial for preserving general knowledge. In contrast, the importance of the q_proj layers is relatively weaker across the model. The L2 norms have been log-transformed to facilitate the comparison of the relative significance of these parameters across layers.
Table 2: Ablation experiments.
@@
Ours 3.38 96.80 3.36 49.13 3.37 36.30
- layer
- layer, element
@@ 6.5 Ablation Study
To investigate the role of different components in our proposed ALoRA, we conducted ablation studies by selectively removing certain structures and observing the resulting impact. Specifically, we excluded two sets of components: (1) no layer importance (eliminating the differentiation of importance among layers), and (2) no layer and element importance (i.e. training $\nu$ task first and subsequently training $\mu$ task, imposing no regularization constraints throughout the process). The slowdown relative to ALoRA after disabling above components are shown in Table 2. We observe that layerwise importance
@@ 7 Conclusion
This paper addresses the critical issue of catastrophic forgetting in large language models (LLMs) during domain-specific fine-tuning. We propose a novel fine-tuning framework to preserve general-purpose knowledge while enabling effective domain-specific learning by utilizing a dual-objective optimization strategy. Our approach ensures that general capabilities are retained, while the model adapts efficiently to new domains, minimizing knowledge degradation in tasks outside the fine-tuned domain. Additionally, we introduce a layer-wise coefficient to adjust the balance between regularization loss and cross-entropy loss dynamically. This adjustment accounts for the varying contributions of different layers to both generalization and domain-specific learning. Extensive experiments in scientific and biological domains show that our framework effectively mitigates catastrophic forgetting and achieves superior performance in domain-specific tasks.

### v2

@@ 5 Experiments
5.1 Backbone LLMs and Baseline Methods
Following the previous work (Xiang et al., 2024), two mainstream LLMs are used for the evaluation of our method: (1) GPT-J (Wang & Komatsuzaki, 2021) is a GPT-2-like causal language model trained on the Pile dataset. It is suitable for various understanding and generation tasks. (2) LLaMA-3 (Dubey et al., 2024) is the third-generation open-source LLM. It is designed with enhanced efficiency and scalability, offering state-of-the-art performance across various benchmarks. These models vary in architecture and parameter count, enabling a robust evaluation of our method.
@@ 5.1 Backbone LLMs and Baseline Methods
5.2 Tasks, Metrics, and Hyperparameters
$\mu$ Task (Domain Ability): The $\mu$ task evaluates the ability to adapt to specific tasks while mitigating catastrophic forgetting of general knowledge. We select three representative tasks: (1) Medical task: MedMCQA dataset (Pal et al., 2022). (2) Scientific task: SciQ dataset (Welbl et al., 2017). (3) Physics task: PiQA dataset (Bisk et al., 2020).
@@ 6 Results and Analysis
6.1 Comparison of General and Domain Capabilities
As shown in Table 1, our method achieves better preservation of general ability (as reflected by the lowest PPL) while maintaining domain-specific accuracy comparable to, or even better than, previous methods. This demonstrates that our approach effectively balances domain accuracy and general perplexity.
@@ 6.2 Complexity Comparision
We compare our HLoRA method with the previous SOTA method, EWCLoRA, from two aspects: the time required for importance calculation and the storage memory needed. As shown in Figure 5, our method is nearly 20 times faster and requires only 10$\sim$15of the storage memory compared to EWCLoRA, demonstrating the practicality of ours.
Time Complexity: The experiments were conducted on an A800 GPU to evaluate the time complexity of our method in comparison with EWCLoRA. For EWCLoRA, the Fisher matrix computation followed the approach described in the original paper, using 20,000 randomly sampled data points from the Pile dataset with a maximum batch size of 8. In contrast, for our method, the time measurement was based on 5 training epochs, a setting determined through empirical evaluation to achieve optimal performance. The experimental results show that for GPT-J-6B and LLaMA-3-3B, EWCLoRA requires 27.17 and 25.97 hours, respectively, to compute the importance matrix, while our HLoRA method only takes 1.15 and 1.19 hours.
@@ 6.3 Regularization Coefficient Analysis
Figure 6 demonstrate the effect of the regularization coefficient $\varphi$ in Equation 16 on PPL and accuracy across three tasks. As $\varphi$ increases, PPL gradually decreases, indicating a stronger emphasis on preserving general ability. Higher values of $\varphi$ correspond to better general ability retention. However, as shown in Figure 6b, increasing $\varphi$ negatively impacts the average accuracy on PiQA. Thus, $e^{-3}$ is selected as the optimal value for the regularization coefficient to balance task performance and general ability (lower PPL).
@@ 6.4 Parameters Importance Visualization
Figure 7 highlights the importance in Equation 13 of q_proj and v_proj layers for general capabilities during the LoRA fine-tuning process. The heatmap illustrates that the v_proj layers, particularly in the first four and the last layer, are crucial for preserving general knowledge. In contrast, the importance of the q_proj layers is relatively weaker across the model. The L2 norms have been log-transformed to facilitate the comparison of the relative significance of these parameters across layers.
Table 2: Ablation experiments. (layer: layer-wise weighted regularization, element: element-wise regularization.)
@@
Ours 3.35 96.80 3.40 50.49 3.34 36.10
- layer 3.36 96.30 3.41 49.95 3.35 35.62
- layer, element 3.39 96.20 3.52 49.51 3.37 33.66
@@ 6.5 Ablation Study
As shown in Table 2, to investigate the role of different components in our proposed HLoRA, we conduct ablation studies by selectively removing certain structures and observing the resulting impact. Specifically, we exclude two sets of components: (1) layer: eliminating the differentiation of importance among layers, and (2) layer, element: removing both layer-wise and element-wise importance, i.e., training the $\nu$ task first and then training the $\mu$ task without imposing any regularization constraints throughout the process. Upon removing the two components, the performance of methods based on two backbone LLMs declines across three datasets, thereby highlighting the effectiveness of the layer-wise and element-wise importance introduced.
@@ 7 Conclusion
This paper addresses the critical issue of catastrophic forgetting in large language models (LLMs) during domain-specific fine-tuning. We propose a novel fine-tuning framework that preserves general capabilities while enabling efficient adaptation to new domains, minimizing knowledge loss in tasks outside the fine-tuned domain. Additionally, we introduce a layer-wise coefficient to adjust the balance between regularization loss and cross-entropy loss dynamically. This adjustment accounts for the varying contributions of different layers to both generalization and domain-specific learning. Extensive experiments in scientific, physical, and medical tasks show that our framework effectively mitigates catastrophic forgetting while maintaining performance in domain-specific tasks.